At the end of the day, mike found that the total cash register receipts at the motel where he works amounted to $2725. This included the 9% sales tax charged. Find the amount of the sales tax.
Answer by rfer(16322)

2725/1.09=2500
2725-2500=225 tax
225/2500==0.09=9%